# Simple Estimation of Semiparametric Models with Measurement Errors

### v3

@@ Abstract
We develop a practical way of addressing the Errors-In-Variables (EIV) problem in the Generalized Method of Moments (GMM) framework. We focus on the settings in which the variability of the EIV is a fraction of that of the mismeasured variables, which is typical for empirical applications. For any initial set of moment conditions our approach provides a “corrected” set of moment conditions that are robust to the EIV. We show that the GMM estimator based on these moments is $\sqrt{n}$-consistent, with the standard tests and confidence intervals providing valid inference. This is true even when the EIV are so large that naive estimators (that ignore the EIV problem) are heavily biased with their confidence intervals having 0% coverage. Our approach involves no nonparametric estimation, which is especially important for applications with many covariates, and settings with multivariate or non-classical EIV. In particular, the approach makes it easy to use instrumental variables to address EIV in nonlinear models.
Keywords: errors-in-variables, nonstandard asymptotic approximation, nonclassical measurement errors, nonparametric identification, instrumental variables
@@ 1 Introduction
Measurement errors are a common problem for empirical studies. Addressing the Errors-In-Variables (EIV) bias in nonlinear models requires elaborate strategies. [^1] See Hausman et al. (1991, 1995); Newey (2001); Schennach (2007); Li (2002); Schennach (2004); Chen et al. (2005); Hu and Schennach (2008); Schennach (2014); Wilhelm (2019), among others. Despite the fundamental theoretical progress in identification and estimation of nonlinear models with EIV, the problem of EIV is still rarely addressed in empirical work outside of linear specifications.
@@
The measurement error can be classical or non-classical. We will refer to $g\left(\cdot\right)$ as the original moment function, since it would have been valid had the researcher observed $X_{i}^{\ast}$. A naive GMM estimator (that ignores the EIV and uses $X_{i}$ in place of $X_{i}^{\ast}$) based on $g\left(\cdot\right)$ is biased because $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]\neq 0$, in contrast to equation (1).
@@
Even in this well-studied example of nonlinear regression, estimation in the presence of the EIV is a difficult problem. Importantly, nonlinear instrumental variable regression estimator cannot be used, since it is inconsistent in the presence of EIV (Amemiya, 1985). The existing approaches typically require nonparametric estimation that can be impractical in many empirical applications. In contrast, in this paper, we develop an alternative class of estimators, that are essentially GMM estimators that modify the original moment functions $g\left(\cdot\right)$ in a way that makes the moment conditions robust to the EIV. In particular, our approach makes it easy to use instrumental variables to address EIV in nonlinear models.
To provide a practical estimation approach for the general class of models (1), we focus on the empirical settings in which the researcher believes the variability of the measurement error to be at most a fraction of the variability of the mismeasured variable, i.e., the noise-to-signal ratio $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ to be moderate, e.g., $\tau\lesssim 0.5$ (we later discuss how the appropriateness of this assumption can be checked). The absolute magnitude of the measurement error $\sigma_{\varepsilon}$ does not need to be small. Focusing on these settings allows us to isolate the most important aspects of the problem and, as result, to develop a simple estimator, which does not require any nonparametric estimation or simulation. Such simple estimation becomes possible because in these settings we can obtain a simple approximation of the EIV bias of the moment conditions as a function of $\theta$.
@@
To make these ideas precise and to study the properties of the proposed estimators, we develop an asymptotic theory using a nonstandard asymptotic approximation that models $\tau$ as slowly shrinking with the sample size. Standard asymptotics considers $\tau$ to be constant, which implies that as $n\rightarrow\infty$ the bias of a naive estimator dwarfs its sampling variability: the bias is constant while the standard errors shrink proportionally to $1/\sqrt{n}$. As a result, under the standard asymptotics, the problem of removing the EIV bias becomes central in the analysis, with relatively little attention paid to the sampling variability of estimators. However, this focus does not seem to be appropriate in many empirical applications, in which the researcher does not expect the potential EIV bias to be several orders of magnitude larger than the standard errors. [^3] Such empirical settings appear to be widespread. Although the concerns about measurement errors are often raised, the majority of applied work does not explicitly correct the EIV bias in nonlinear models, and instead implicitly or explicitly argues or conjectures that the EIV bias is likely not to be too large. See also the review of Bound et al. (2001). By considering $\tau$ as drifting towards zero with the sample size, our approach provides a better guidance on construction of EIV robust estimators with good finite sample properties when $\tau$ is small or moderate. [^4] Nonstandard asymptotic approximations with drifting parameters are often used to obtain better approximations of the finite sample behavior of estimators and tests. For example, in the instrumental variable regression settings, to consider the settings with relatively small first stage coefficients, Staiger and Stock (1997) model them as shrinking with $n$. It is important to keep in mind that such nonstandard asymptotic approximations are merely mathematical tools. One should not take them literally and think of parameters somehow changing if more data is collected.
@@
The MERM estimator is structurally different from the existing approaches that require nonparametric estimation of some nuisance parameters, for example, of the density $f_{X_{i}^{\ast}|W}$. Avoiding nonparametric estimation has at least two advantages. First, since the majority of empirical applications include at least a handful of additional covariates $W_{i}$, nonparametric estimation is often infeasible due to the curse of dimensionality. Because the MERM estimator does not involve any nonparametric estimation, it can be used in applications with a relatively large number of additional covariates $W_{i}$, and remains feasible even in the more complicated settings, including multi-equation and structural models, and applications with multiple mismeasured variables $X_{i}$. Second, estimation of infinite-dimensional nuisance parameters is typically more demanding towards the sources of identification available in the data, for example, requiring an instrumental variable with a large support (continuously distributed). In contrast, having a discrete instrument is sufficient for the MERM approach because the nuisance parameter $\gamma_{0}$ is finite-dimensional.
@@
Related Literature Chen et al. (2011), Schennach (2016), and Schennach (2020) provide excellent overviews of the measurement error literature.
The existing semiparametric approaches to estimation and inference in models with EIV involve nonparametric estimation of infinite-dimensional nuisance parameters (e.g., Chesher2000WP,Li2002JoE,Schennach2004Ecta,Schennach2007Ecta,HuSchennach2008Ecta,SchennachHu2013JASA,Song2015JoE), simulation (e.g., Schennach2014Ecta-ELVIS), or both (e.g., Newey2001REStat,WangHsiao2011JoE)))). The exceptions include models with linear and polynomial regression functions (see HINP1991JoE,HausmanNeweyPowell1995JoE)))), and Gaussian control variable models such as Probit and Tobit with endogeneity (see SmithBlundell1986Ecta,RiversVuong1988JoE).
@@
We are able to provide such an estimator because we focus on the models with moderate measurement errors. Modeling the variance of the measurement error as shrinking to zero with the sample size is a popular approach in Statistics. The method has been proposed by Wolter and Fuller (1982), who used it to construct an approximate MLE estimator of a nonlinear regression model with Gaussian errors. Following their approach, the Statistics literature has mainly focused on the settings where the moments of the EIV needed to bias correct the estimators are either known or can be directly estimated from the available data such as repeated measurements (e.g., CarrollStefanski1990JASA,CarrollEtAl2006Book-ME). In Economics, such data are relatively rare. The use of approximations with shrinking variance of measurement errors in Econometrics literature has been pioneered by Kadane (1971), Amemiya (1985), and Chesher (1991). Such approximations have been used to check the sensitivity of naive estimators to the EIV by considering how the estimates change as the unknown moments of the measurement errors vary within some set of plausible values, e.g., see Chesher and Schluter (2002), Chesher et al. (2002), Battistin and Chesher (2014), Chesher (2017), and Hong and Tamer (2003).
@@
This paper differs from the earlier literature in several ways. First, it presents a way to estimate the unknown nuisance parameters (moments of the measurement errors) jointly with the parameters of interest. As a result, the approach can, for example, use instrumental variables as a source of identification. Second, the method applies to a very general class of semiparametric models specified by moment conditions. Third, the MERM approach allows the measurement errors to have larger magnitudes than most of the papers in the earlier literature; this is achieved by the MERM approach recursively bias correcting the bias correction terms. Fourth, our approach allows relaxing the assumption of classical measurement errors.
@@
The most widespread approach to identification of the EIV models in economic applications is to use instrumental variables, e.g., see Hausman et al. (1991); Newey (2001); Schennach (2007); Wang and Hsiao (2011). In a recent paper, Hahn et al. (2021) reconsider the regression model in Amemiya (1990) using a bias correction similar to ours. When proper excluded variables are not available, researchers have considered using higher moments of $X_{i}$ as instruments, e.g., see Reiersøl (1950); Lewbel (1997); Erickson and Whited (2002); Schennach and Hu (2013); Ben-Moshe et al. (2017). When available, repeated measurements can also be used to identify the model, e.g., see Hausman et al. (1991); Li and Vuong (1998); Li (2002); Schennach (2004). The MERM estimator accommodates these identification approaches within a unified estimation framework.
@@
The power of the general MERM approach can be illustrated in the NLR model. For example, when a candidate instrumental variable is available, the conditions it needs to satisfy are much weaker than what is required by many existing approaches. Availability of a discrete instrument is sufficient for identification; and the instrument is allowed to have heterogeneous impact on covariates $X_{i}^{\ast}$. One can also take a nonclassical, nonlinear (e.g., discretized or censored), or biased measurement of $X_{i}^{\ast}$ as an instrument in the MERM approach. We discuss identification in Section 2.4. In addition, in a related paper Evdokimov and Zeleneev (2022) study nonparametric regression with EIV using the $\tau\rightarrow 0$ approximation, and demonstrate that the MERM approach can also be motivated from a nonparametric perspective.
@@
Kitamura et al. (2013); Andrews et al. (2017); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), among others, develop tools for estimation and inference in GMM, which are robust to general perturbation or misspecification of the true data generating process. They focus on the settings in which these perturbations are sufficiently small, so that naive estimators remain $\sqrt{n}$-consistent, and their biases are of the same order of magnitude as their standard errors. In contrast, we focus on more specific forms of data contamination due to the EIV. This allows the MERM approach to remain valid even in the settings with larger measurement errors, in which naive estimators may have slower than $\sqrt{n}$ rates of convergence.
@@
The MERM approach also provides a useful foundation for dealing with EIV in more complicated settings. Evdokimov and Zeleneev (2018) utilize the MERM framework to address an issue of nonstandard inference, which turns out to arise generally when EIV models are identified using instrumental variables. Evdokimov and Zeleneev (2019) extend the analysis of this paper to long panel and network settings.
@@
Organization of the paper Section 2 introduces the Moderate Measurement Error framework and the proposed MERM estimator. Section 3 presents several Monte Carlo experiments that illustrate finite sample properties of the MERM estimators. Section 4 considers several extensions of the framework.
@@ 2 Moderate Measurement Errors Framework
To present the main ideas we first consider the case of univariate $X_{i}^{\ast}$ and classical measurement error $\varepsilon_{i}$. Later we consider multivariate $X_{i}^{\ast}$ and non-classical measurement errors $\varepsilon_{i}$.
To develop a practical estimation approach for general moment condition models we focus on the settings in which $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ is small or moderate. We consider an asymptotic approximation with $\tau_{n}\equiv\tau\rightarrow 0$ as $n\rightarrow\infty$. Note that economically meaningful parameters are usually invariant to rescaling of $X_{i}^{\ast}$. Likewise, the extent of the EIV problem does not change with such rescaling.
[…]

### v4

@@ Abstract
We develop a practical way of addressing the Errors-In-Variables (EIV) problem in the Generalized Method of Moments (GMM) framework. We focus on the settings in which the variability of the EIV is a fraction of that of the mismeasured variables, which is typical for empirical applications. For any initial set of moment conditions our approach provides a “corrected” set of moment conditions that are robust to the EIV. We show that the GMM estimator based on these moments is $\sqrt{n}$-consistent, with the standard tests and confidence intervals providing valid inference. This is true even when the EIV are so large that naive estimators (that ignore the EIV problem) are heavily biased with their confidence intervals having 0% coverage. Our approach involves no nonparametric estimation, which is especially important for applications with many covariates and settings with multivariate EIV. In particular, the approach makes it easy to use instrumental variables to address EIV in nonlinear models.
Keywords: errors-in-variables, nonstandard asymptotic approximation, nonparametric identification, instrumental variables
@@ 1 Introduction
Measurement errors are a common problem for empirical studies. Addressing the Errors-In-Variables (EIV) bias in nonlinear models requires elaborate strategies. [^1] See HINP1991JoE; HausmanNeweyPowell1995JoE; Newey2001REStat; Schennach2007Ecta; Li2002JoE; Schennach2004Ecta; ChenHongTamer2005ReStud; HuSchennach2008Ecta; Schennach2014Ecta-ELVIS; Wilhelm2019WP-TestingForME, among others. Despite the fundamental theoretical progress in identification and estimation of nonlinear models with EIV, the problem of EIV is still rarely addressed in empirical work outside of linear specifications.
@@
We will refer to $g\left(\cdot\right)$ as the original moment function, since it would have been valid had the researcher observed $X_{i}^{\ast}$. A naive GMM estimator (that ignores the EIV and uses $X_{i}$ in place of $X_{i}^{\ast}$) based on $g\left(\cdot\right)$ is biased because $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]\neq 0$, in contrast to equation (1).
@@
Even in this well-studied example of nonlinear regression, estimation in the presence of the EIV is a difficult problem. Importantly, nonlinear instrumental variable regression estimator cannot be used, since it is inconsistent in the presence of EIV (Amemiya1985JoE). The existing approaches typically require nonparametric estimation that can be impractical in many empirical applications. In contrast, in this paper, we develop an alternative class of estimators, that are essentially GMM estimators that modify the original moment functions $g\left(\cdot\right)$ in a way that makes the moment conditions robust to the EIV. In particular, our approach makes it easy to use instrumental variables to address EIV in nonlinear models.
To provide a practical estimation approach for the general class of models (1), we focus on empirical settings in which the researcher believes the variability of the measurement error to be at most a fraction of the variability of the mismeasured variable, i.e., the noise-to-signal ratio $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ to be moderate. The absolute magnitude of the measurement error $\sigma_{\varepsilon}$ does not need to be small. Existing validation studies provide insights into the magnitude of $\tau$ for some key economic variables and datasets. BoundKrueger1991JoLaborEcon consider log-earnings in the Current Population Survey (CPS) data matched to the Social Security payroll records. Their estimates of the variance of the measurement errors correspond to $\tau$ of approximately $0.47$ and $0.30$ for the subsamples of men and women, respectively. BoundBrownDuncanRodgers1994JoLaborEcon consider a validation study of Panel Study of Income Dynamics (PSID). Their estimates imply $\tau=0.39-0.66$ for log-earnings and $\tau=0.63-0.76$ for hours worked. Pischke1995JBES estimates correspond to $\tau=0.39-0.50$ for the log-earnings in PSID. AshenfelterKrueger1994AER assess the mismeasurement in the years of education; their estimates correspond to $\tau=0.30-0.37$.
Focusing on these settings allows us to isolate the most important aspects of the problem and, as result, to develop a simple estimator, which does not require any nonparametric estimation or simulation. Such simple estimation becomes possible because in these settings we can obtain a simple approximation of the EIV bias of the moment conditions as a function of $\theta$.
@@
To make these ideas precise and to study the properties of the proposed estimators, we develop an asymptotic theory using a nonstandard asymptotic approximation that models $\tau$ as slowly shrinking with the sample size. Standard asymptotics considers $\tau$ to be constant, which implies that as $n\rightarrow\infty$ the bias of a naive estimator dwarfs its sampling variability: the bias is constant while the standard errors shrink proportionally to $1/\sqrt{n}$. As a result, under the standard asymptotics, the problem of removing the EIV bias becomes central in the analysis, with relatively little attention paid to the sampling variability of estimators. However, this focus does not seem to be appropriate in many empirical applications, in which the researcher does not expect the potential EIV bias to be several orders of magnitude larger than the standard errors. [^3] Such empirical settings appear to be widespread. Although the concerns about measurement errors are often raised, the majority of applied work does not explicitly correct the EIV bias in nonlinear models, and instead implicitly or explicitly argues or conjectures that the EIV bias is likely not to be too large. By considering $\tau$ as drifting towards zero with the sample size, our approach provides a better guidance on construction of EIV robust estimators with good finite sample properties when $\tau$ is small or moderate. [^4] Nonstandard asymptotic approximations with drifting parameters are often used to obtain better approximations of the finite sample behavior of estimators and tests. For example, in the instrumental variable regression settings, to consider the settings with relatively small first stage coefficients, StaigerStock1997 model them as shrinking with $n$. It is important to keep in mind that such nonstandard asymptotic approximations are merely mathematical tools. One should not take them literally and think of parameters somehow changing if more data is collected.
@@
The MERM estimator is structurally different from the existing approaches that require nonparametric estimation of some nuisance parameters, for example, of the density $f_{X^{\ast}|Z,W}$. Avoiding nonparametric estimation has at least two advantages. First, since the majority of empirical applications include additional covariates $W_{i}$, nonparametric estimation is often infeasible due to the curse of dimensionality. Because the MERM estimator does not involve any nonparametric estimation, it can be used in applications with a relatively large number of additional covariates $W_{i}$, and remains feasible even in the more complicated settings, including multi-equation and structural models, and applications with multiple mismeasured variables $X_{i}$. Second, estimation of infinite-dimensional nuisance parameters is typically more demanding towards the sources of identification available in the data, for example, requiring an instrumental variable with a large support (continuously distributed). In contrast, having a discrete instrument is sufficient for the MERM approach because the nuisance parameter $\gamma_{0}$ is finite-dimensional.
For example, in Section 3.3 we consider estimation of the model of multinomial choice among three modes of transportation. A leading alternative approach to the errors-in-variables problem in this model is the semi-nonparametric sieve-MLE estimator advocated by HuSchennach2008Ecta; CarrollChenHu2010JoNS, among others. This approach requires, among other things, estimating the conditional density $f_{X^{\ast}|Z,W}$ of $X_{i}^{\ast}$ given the instrument $Z_{i}$ and covariates $W_{i}$. In this empirical example, $W_{i}$ includes four continuously distributed covariates (two continuously distributed characteristics per choice) and a discrete one, while scalar $X_{i}^{\ast}$ and $Z_{i}$ are also continuous. Thus, $f_{X^{\ast}|Z,W}$ is a function of six continuous and one discrete variable. Hence, for typical sample sizes, estimating $f_{X^{\ast}|Z,W}$ in this example is infeasible due to the curse of dimensionality. In contrast, as the results of Section 3.3 demonstrate, the MERM approach is practical and effective in this application, in part because it avoids estimation of the high-dimensional nuisance functions like $f_{X^{\ast}|Z,W}$ altogether.
@@
Related Literature ChenHongNekipelov2011JEL, Schennach2016AnnRev, and Schennach2020HB-ME provide excellent overviews of the measurement error literature.
The existing semiparametric approaches to estimation and inference in models with EIV involve nonparametric estimation of infinite-dimensional nuisance parameters (e.g., Chesher2000WP,Li2002JoE,Schennach2004Ecta,Schennach2007Ecta,HuSchennach2008Ecta,SchennachHu2013JASA,Song2015JoE), simulation (e.g., Schennach2014Ecta-ELVIS), or both (e.g., Newey2001REStat,WangHsiao2011JoE)). The exceptions include models with linear and polynomial regression functions (see HINP1991JoE,HausmanNeweyPowell1995JoE)), and Gaussian control variable models such as Probit and Tobit with endogeneity (see SmithBlundell1986Ecta,RiversVuong1988JoE).
@@
We are able to provide such an estimator because we focus on the models with moderate measurement errors. Modeling the variance of the measurement error as shrinking to zero with the sample size is a popular approach in Statistics. The method has been proposed by WolterFuller1982AS, who used it to construct an approximate MLE estimator of a nonlinear regression model with Gaussian errors. Following their approach, the Statistics literature has mainly focused on the settings where the moments of the EIV needed to bias correct the estimators are either known or can be directly estimated from the available data such as repeated measurements (e.g., CarrollStefanski1990JASA,CarrollEtAl2006Book-ME). In Economics, such data are relatively rare. The use of approximations with shrinking variance of measurement errors in Econometrics literature has been pioneered by Kadane1971Ecta, Amemiya1985JoE, and Chesher1991Biomet. Such approximations have been used to check the sensitivity of naive estimators to the EIV by considering how the estimates change as the unknown moments of the measurement errors vary within some set of plausible values, e.g., see ChesherSchluter2002ReStud, ChesherDumanganeSmith2002JoE, BattistinChesher2014JoE, Chesher2017JoE, and HongTamer2003JoE. BoundBrownMathiowetz2001HBoE review a broad list of validation studies matching standard economic dataset to administrative records. The estimates they report suggest that the measurement errors of moderate magnitude are typical for empirical applications. This suggests that the approach developed in this paper could prove valuable for a wide range of applied work.
@@
This paper differs from the earlier literature in several ways. First, it presents a way to estimate the unknown nuisance parameters (moments of the measurement errors) jointly with the parameters of interest. As a result, the approach can, for example, use instrumental variables as a source of identification. Second, the method applies to a very general class of semiparametric models specified by moment conditions. Third, the MERM approach allows the measurement errors to have larger magnitudes than most of the papers in the earlier literature; this is achieved by the MERM approach recursively bias correcting the bias correction terms.
@@
The most widespread approach to identification of the EIV models in economic applications is to use instrumental variables, e.g., see HINP1991JoE; Newey2001REStat; Schennach2007Ecta; WangHsiao2011JoE. In a recent paper, HahnHausmanKim2021EL reconsider the regression model in Amemiya1990JoE using a bias correction similar to ours. When proper excluded variables are not available, researchers have considered using higher moments of $X_{i}$ as instruments, e.g., see Reiersol1950Ecta; Lewbel1997Ecta; EricksonWhited2002ET; SchennachHu2013JASA; BenMosheDHaultfeuilleLewbel2017JoE. When available, repeated measurements can also be used to identify the model, e.g., see HINP1991JoE; LiVuong1998JoE; Li2002JoE; Schennach2004Ecta. The MERM estimator accommodates these identification approaches within a unified estimation framework.
@@
The power of the general MERM approach can be illustrated in the NLR model. For example, when a candidate instrumental variable is available, the conditions it needs to satisfy are much weaker than what is required by many existing approaches. Availability of a discrete instrument is sufficient for identification; and the instrument is allowed to have heterogeneous impact on covariates $X_{i}^{\ast}$. One can also take a nonclassical, nonlinear (e.g., discretized or censored), or biased measurement of $X_{i}^{\ast}$ as an instrument in the MERM approach. We discuss identification in Section 2.4. In addition, in a related paper EvdokimovZeleneev2022WP-NPID study nonparametric regression with EIV using the $\tau\rightarrow 0$ approximation, and demonstrate that the MERM approach can also be motivated from a nonparametric perspective.
@@
KitamuraOtsuEvdokimov2013Ecta; AndrewsGentzkowShapiro2017QJE; ArmstrongKolesar2021QE; BonhommeWeidner2021QE, among others, develop tools for estimation and inference in GMM, which are robust to general perturbation or misspecification of the true data generating process. They focus on the settings in which these perturbations are sufficiently small, so that naive estimators remain $\sqrt{n}$-consistent, and their biases are of the same order of magnitude as their standard errors. In contrast, we focus on more specific forms of data contamination due to the EIV. This allows the MERM approach to remain valid even in the settings with larger measurement errors, in which naive estimators may have slower than $\sqrt{n}$ rates of convergence.
@@
The MERM approach also provides a useful foundation for dealing with EIV in more complicated settings. EvdokimovZeleneev2018WP-Inference utilize the MERM framework to address an issue of nonstandard inference, which turns out to arise generally when EIV models are identified using instrumental variables. EvdokimovZeleneev2019WP-Panel extend the analysis of this paper to long panel and network settings.
@@
Organization of the paper Section 2 introduces the Moderate Measurement Error framework and the proposed MERM estimator. Section 3 presents several Monte Carlo experiments that illustrate finite sample properties of the MERM estimators. Section 4 considers several extensions of the framework. A supplementary appendix contains all proofs and additional results for the numerical and empirical illustrations.
@@ 2 Moderate Measurement Errors Framework
To present the main ideas we first consider the case of univariate $X_{i}^{\ast}$. We will consider multivariate $X_{i}^{\ast}$ later. We assume that the measurement error is classical, i.e., that $\varepsilon_{i}$ is independent of $X_{i}^{*}$ and $S_{i}$; later we will discuss how this assumption can be relaxed. Following the rest of the literature, we assume that $\mathbb{E}\left[\varepsilon_{i}\right]=0$. [^5] A location normalization such as $\mathbb{E}\left[\varepsilon_{i}\right]=0$ is usually necessary because it is not possible to separately identify the means $\mathbb{E}\left[X_{i}^{\ast}\right]$ and $\mathbb{E}\left[\varepsilon_{i}\right]$.
To develop a practical estimation approach for general moment condition models we focus on the settings in which $\tau\equiv\sigma_{\varepsilon}/\sigma_{X^{\ast}}$ is small or moderate. We consider an asymptotic approximation with $\tau_{n}\equiv\tau\rightarrow 0$ as $n\rightarrow\infty$.
Note that economically meaningful parameters are usually invariant to rescaling of $X_{i}^{\ast}$. Likewise, the extent of the EIV problem does not change with such rescaling. For simplicity of exposition, it is convenient to assume that $X_{i}^{\ast}$ is scaled so that $\sigma_{X^{\ast}}$ is of order one and, correspondingly, the moments $\mathbb{E}[\left|\varepsilon_{i}\right|^{k}]\propto\tau_{n}^{k}$ decrease with $k$ when $\tau_{n}<1$. For example, this could be ensured by normalizing observed $X_{i}$ to have $\sigma_{X}=1$. Let us stress that this normalization is used only to simplify the exposition; as we show in Appendix G, the proposed MERM estimator does not require any normalizations in practice.